# EDEN: Efficient Dual-Layer Exploration Planning for Fast UAV Autonomous Exploration in Large 3-D Environments

### v1

@@ Abstract
Efficient autonomous exploration in large-scale environments remains challenging due to the high planning computational cost and low-speed maneuvers. In this paper, we propose a fast and computation-efficient dual-layer exploration planning method. The insight of our dual-layer method is efficiently finding an acceptable long-term region routing and greedily exploring the target in the region of the first routing area with high speed. Specifically, the proposed method finds the long-term area routing through an approximate algorithm to ensure real-time planning in large-scale environments. Then, the viewpoint in the first routing region with the lowest curvature-penalized cost which can effectively reduce decelerations caused by sharp turn motions will be chosen as the next exploration target. To further speed up the exploration, we adopt an aggressive and safe exploration-oriented trajectory to enhance exploration continuity. The proposed method is compared to state-of-the-art methods in challenging simulation environments. The results show the proposed method outperforms other methods in terms of exploration efficiency, computational cost, and trajectory speed. We also conduct real-world experiments to validate the effectiveness of the proposed method. The code will be open-sourced [^1] https://github.com/NKU-MobFly-Robotics/EDEN.
Index Terms:
Autonomous exploration, motion planning, trajectory planning.
@@ I Introduction
Umanned aerial vehicles (UAVs) are popular in many 3-D exploration relevant applications, e.g., search and rescue [1, 2], environment inspection [3, 4], environment reconstruction [5, 6, 7, 8], among others. However, most existing methods remain low efficiency, low trajectory speed, and high computational cost.
@@
Existing exploration methods can be coarsely separated into two kinds of methods: greedy methods and optimal-routing-based methods. Greedy methods attempt to select the target that maximizes the reward in a short period. The original frontier-based method [9] selects the closest frontier, i.e., the boundary voxel of explored and unknown space as an exploration target. To improve the UAV speed, [10] propose to select a frontier in camera field-of-view (FOV) to minimize the velocity change. By combining Next-Best-Views and rapid random tree (RRT), receding-horizon-next-best-views (RH-NBVP) [11] is proposed, and its variant methods [12, 13, 14, 15] show high exploration efficiency in small environments. However, these greedy methods suffer from massive back-and-forth maneuvers in large-scale environments due to the lack of long-term consideration. Another type of method, optimal-routing-based methods [16, 17, 18, 19, 20, 21], make decisions by constructing and solving routing problems to minimize the total routing length. The work in [21] proposes to find a coverage path by solving the traveling salesman problem (TSP). The method proposed by [16] constructs and solves the asymmetrical traveling salesman problem (ATSP) to find the shortest routing of frontiers. To further reduce back-and-forth maneuvers, [18] increases the routing priority of isolated frontiers. By solving routing problems, these methods dramatically reduce back-and-forth maneuvers. However, the computational cost grows fast with the number of routing nodes, which makes these methods have to execute stops for planning in large-scale environments. What is more, the work in [22] shows that the speed of UAVs in exploration still has great potential to be stimulated. Therefore, a computation-efficient, far-sighted, and high-speed method is expected.
@@
Motivated by the above drawbacks of existing methods, we propose an Efficient Dual-layer Exploration plaNning method (EDEN) which can plan efficient exploration motions in hundred-meter-level 3-D environments in real-time. Different from greedy methods and optimal-routing-based methods, the proposed dual-layer method first finds a long-term sub-optimal exploration region routing efficiently and then decides the short-term targets that minimizes the exploration motion cost greedily. Specifically, we improve the Christofides algorithm, which is an approximate algorithm for solving TSP, to an exploration-oriented heuristic Christofides (EOHC) algorithm to find a sub-optimal but satisfying tour for ATSP with low computational cost. Then, a viewpoint for safety and two viewpoints with the highest curvature-penalized exploration gain in the first exploration region of the routing will be selected for high-speed trajectory planning. Then, inspired by [23], we propose an aggressive and safe exploration-oriented (ASEO) trajectory for high-speed continuous exploration. The ASEO trajectory is composed of three parts: an exploring part, a continuous part, and a safety part. The time of the exploring part, the time of the continuous part, and the feasibility of the safety part are optimized together to find an aggresive trajectory and a backup trajectory for safety.
@@
We evaluate the proposed method by comparing it against state-of-the-art methods in various challenging environments. The results show that the proposed method has $5.2\%-14.0\%$ higher exploration efficiency, $10.4\%-15.5\%$ faster UAV speed, and $28.9\%-74.5\%$ lower computational cost. Further, the proposed method is deployed on a UAV and is tested in real-world environments to evaluate its effectiveness. The contributions of this work are summarized as follows:
@@
    We propose an exploration-oriented heuristic Christofides (EOHC) algorithm for efficiently solving large-scale ATSP. The proposed EOHC can find a satisfactory region routing with a low computational cost. The real-time performance of EOHC can be guaranteed even in hundred-meter-level environments.
@@
    To take advantage of the current UAV speed and enhance exploration continuity, we propose a heuristic curvature-penalized continuous gain for the short-term exploration targets choosing.
@@
    To make the UAV keep high-speed continuous exploration in unknown environments, we propose an ASEO trajectory planning method that ensures trajectory efficiency and safety at the same time.
@@
    Massive comparative simulations are conducted to evaluate the proposed method. We also test the proposed method by real-world experiments. We will release our source code.
@@ II-A Greedy Exploration Planning
The frontier-based method is first proposed by Yamauchi et al. [9]. It selects the frontier voxel that is closest to the robot as the exploration target. To minimize the UAV velocity change, [10] selects the frontier voxel that is located at the boundary of camera FOV as the exploration target. Different from the greedy frontier-based methods that only consider the next step exploration, Bircher et al. [11] propose RH-NBVP which samples a set of exploration motion sequences and executes the most informative sequence’s first step and it shows high efficiency in short-term exploration. Nevertheless, in large-scale environments, it suffers from sample failure. To cope with this problem, Selin et al. [24] use frontier to guide sampling when a local sample fails. However, long-distance sampling is time-consuming and it does not guarantee success even with the guidance of frontier. To reduce the planning cost and improve long-distance exploration, roadgraphs, which can compress the complex explorable space information and tarversable space information into compact nodes and edges, are widely used in recent proposed greedy methods [12, 25, 14, 26, 27, 28, 29, 30]. Consequently, roadgraphs enable greedy methods of efficient planning in large environments. Although these methods show high efficiency in short-term exploration and low computational cost, they suffer from back-and-forth maneuvers due to a lack of long-term consideration.
@@ II-B Optimal-Routing-Based Exploration Planning
Optimal-routing-based methods usually formulate the exploration target decision problems into routing problems. Optimal-tour-based methods can be roughly devided into frontier-ATSP methods [16, 17, 31, 18] and coverage-path-ATSP methods [32, 20, 21, 5, 19]. Frontier-ATSP methods formulate ATSP using frontiers and find the shortest tour between frontiers. Coverage-path-ATSP methods formulate ATSP with unexplored space and find the shortest coverage path. The main drawback of the optimal-routing-based methods is that they have to spend luxury computation on solving ATSP in large-scale environments. Moreover, counterintuitively, recent works [28, 33] report that optimal-tour-based methods usually do not have an obvious advantage in minimizing the exploration path distance compared to greedy methods. This is because optimal-routing-based methods have implicit optimal assumptions, where frontier-ATSP methods assume a new frontier will not be generated after exploring an existing frontier, while coverage-path-ATSP methods assume all the unexplored space is traversable. Obviously, the above assumptions do not hold in most cases. Therefore, spending too much computation on finding a high-quality ATSP solution is not cost-effective.
II-C Exploration-Oriented Trajectory Planning
Exploration trajectory plays a crucial role in exploration tasks. The early proposed exploration methods [11, 9] stop-and-go motions are used. To utilize the good maneuverabilities of UAVs, Dharmadhikari et, al. [15] adopts motion primitives to speed up the exploration. To further improve the trajectory speed, Zhou et al. [16] optimizes minimum-time trajectory with optimizable terminal velocity and acceleration for high-speed exploration. Although minimum-time trajectory is very efficient, the optimizable terminal velocity and acceleration do not ensure being beneficial to continuously exploring the newly generated frontier after exploring the previous frontier. What is more, the optimizable terminal velocity and acceleration can lead to collision.
@@ II-C Exploration-Oriented Trajectory Planning
The proposed method is composed of a dual-layer exploration target decision module that can plan both the long-term region routing and the short-term viewpoint target, and a trajectory planning module that can plan high-speed continuous exploration trajectories as shown in Fig. 1. A dynamic voxel map with resolution $r_{v}$ is updated by integrating depth image and odometry information. Different from the normal voxel map, the dynamic voxel map only maintains a local part of voxels around the UAV while other voxels are dynamically read and written to save memory usage in large environments. Based on the voxel map, a dynamic topological graph (DTG), is maintained to manage exploration-related information. The traversable information is maintained by DTG’s history nodes and edges between them, and the explorable space information is maintained by DTG’s explorable region of interest (EROI), which is a set of uniform scale cubic regions that can be explored by visiting corresponding active viewpoints. For detailed information, we refer readers to our previous work [34]. The exploration planning utilizes DTG to efficiently formulate ATSP, and ATSP is solved through EOHC to find satisfactory exploration region routing with small computation. Then, three viewpoints, including an exploring viewpoint, a continuous viewpoint, and a safety viewpoint in the first region of the region routing are found for the following trajectory planning. Then, an aggressive and safe exploration-oriented trajectory which separates at the exploring viewpoint, and its two branches end at a continuous viewpoint and a safe viewpoint correspondingly, is optimized. The exploring part of the aggressive and safe exploration-oriented trajectory will be executed.
@@ IV Dual-layer exploration viewpoints planning
In this section, we first propose the EOHC algorithm for finding a satisfactory region routings efficiently in large-scale environments. Then, we use curvature-penalized continuous gain to decide the targets in the first region of the tour for the following high-speed ASEO trajectory planning.
@@
Figure 2: An instance of EOHC-based routing planning, the numbers indicate the lengths of edges. (a) An MST that connects all the explorable history nodes of DTG is extracted. (b) The leaf point, whose distance is the longest to the root node, and its corresponding branch (green) will be executed last. (c) Once, there are multiple siblings for the current node, the sibling (orange dashed curve) whose first node has the shortest distance to the current node will be chosen as the next branch. (d) The final routing found by the proposed EOHC-based routing planning.
IV-A EOHC-Based Global Routing Planning
@@
As analyzed in Section. II.B, the optimal routing of frontiers or unexplored space is very likely not the optimal routing for exploration due to the implicit optimal assumption not being met. Therefore, instead of finding a high-quality of all precise explorable frontiers, the proposed method aims to spend low computation on finding a satisfactory coarse routing of explorable regions. The process of EOHC-based routing planning is shown in Fig. 2.
@@
Firstly, we extract all the history nodes of DTG that connect at least one EROI as the explorable regions. To reduce the computation when formulating ATSP, the distances between all explorable regions and the distances between the robot and all explorable regions are computed by conducting the Dijkstra algorithm on DTG. Thanks to the sparse DTG, the ATSP can be formulated efficiently.
@@
To solve ATSP efficiently, we design an EOHC algorithm, which is improved from the Christofides algorithm [35] and tailored to ATSP. The key idea of the Christofides algorithm is to construct a minimum spanning tree (MST, Fig. 2(a)) using the Prim algorithm and find a walk that goes around the MST. The solution routing of the Christofides algorithm ensures no longer than twice the optimal routing. However, the Christofides algorithm is designed for symmetric TSP. For exploration-oriented ATSP, the robot does not have to return to the original node as stated by [16]. Therefore, to minimize the routing length, the leaf node that is the farthest from the robot and nodes on its corresponding branch (in Fig. 2(b), the green nodes) will be executed last to save the highest cost of returning to the original node. Another problem is how to decide the sequence of other branches. Since all the regions will be explored finally, the proposed method chooses the next node closest to the current node heuristically when the current node has multiple siblings as shown in Fig. 2(c). Finally, a routing for ATSP will be found as shown in Fig. 2(d).
For the worst case, the path length $l_{eohc}$ of the routing solved by the EOHC algorithm holds:
$l_{eohc}\leq 2l^{*}+l_{m},$ (1)
@@
where $l^{*}$ is the optimal solution of the corresponding ATSP, and $l_{m}$ is the longest distance among all the distances between the robot and all the explorable history nodes. The computational complexity of the proposed EOHC is $\mathcal{O}(n^{3})$, where $n$ is the total number of nodes.

### v2

@@ Abstract
Efficient autonomous exploration in large-scale environments remains challenging due to high planning computational cost and low-speed maneuvers. In this paper, we propose an Efficient Dual-layer Exploration plaNning method (EDEN). The insight of our dual-layer planning method is efficiently finding an acceptable long-term region routing and greedily exploring the target in the first region of routing at high speed. Specifically, we propose a long-term region routing approximate algorithm, called “exploration-oriented heuristic double-tree (EOHDT) algorithm”, to ensure real-time planning in large-scale environments. Then, the viewpoint in the first routing region with the highest curvature-penalized score, which can effectively reduce decelerations caused by sharp turn motions, will be chosen as the next exploration target. To further speed up the exploration, we propose an aggressive and safe exploration-oriented trajectory planning approach to enhance exploration continuity and speed. The proposed method is compared with state-of-the-art methods in challenging simulation environments. The results show that the proposed method outperforms other methods in terms of exploration efficiency, computational cost, and trajectory speed. We also conduct real-world experiments to validate the effectiveness of the proposed method. The code will be open-sourced [^1] https://github.com/NKU-MobFly-Robotics/EDEN.
@@ I INTRODUCTION
Unmanned aerial vehicles (UAVs) are popular in many 3-D exploration-relevant applications, e.g., search and rescue [1, 2], environment inspection [3, 4, 5], environment reconstruction [6, 7, 8, 9], healthcare services [10, 11], among others. However, most existing methods suffer from low efficiency, low-speed trajectories, and high computational cost.
@@
Existing exploration methods can be coarsely separated into three kinds of methods: greedy methods, optimal-routing-based methods, and learning-based methods. Greedy methods attempt to select the target that maximizes the reward in a short period. The original frontier-based method [12] selects the closest frontier, i.e., the boundary voxel of explored and unknown space as the exploration target. To improve the UAV speed, the work in [13] proposes to select a frontier in the camera field-of-view (FOV) to minimize the velocity change. By combining next-best-views and rapid random tree (RRT), receding-horizon-next-best-views (RH-NBVP) [14] is proposed, and its variant methods [15, 16, 17, 18] show high exploration efficiency in small environments. However, these greedy methods suffer from massive back-and-forth maneuvers in large-scale environments due to the lack of long-term consideration. The optimal-routing-based methods [19, 20, 21, 22, 23], make decisions by constructing and solving routing problems to minimize the total routing length. The work in [23] proposes to find a coverage path by solving the traveling salesman problem (TSP). The method proposed by [19] constructs and solves the asymmetrical traveling salesman problem (ATSP) to find the shortest routing of frontiers. To further reduce back-and-forth maneuvers, the approach in [20] increases the routing priority of isolated frontiers. By solving routing problems, these methods dramatically reduce back-and-forth maneuvers. However, the computational cost grows fast with the number of routing nodes, which makes these methods have to execute stops to wait for planning results in large-scale environments. In addition, the works in [24, 22] show that the speed of UAVs in exploration has great potential to be stimulated and can be significantly increased, while most existing state-of-the-art methods fail. Some learning-based methods [25, 26, 27] are proposed to explore the large and complex environments. However, they share the common limitation of suboptimal path planning due to discrete action spaces. Therefore, a computation-efficient, far-sighted, and high-speed exploration method is highly expected.
@@
Motivated by the above drawbacks of existing methods, we propose an Efficient Dual-layer Exploration plaNning method (EDEN) which can plan efficient exploration motions in hundred-meter-level 3-D environments in real-time. Different from greedy methods and optimal-routing-based methods, the proposed dual-layer method first finds a long-term sub-optimal region routing efficiently and then decides the short-term targets that can fully take advantage of the current UAV speed. Specifically, we propose the exploration-oriented heuristic double-tree (EOHDT) algorithm to find a sub-optimal satisfactory routing for ATSP with low computational cost. Then, a viewpoint for safety and two viewpoints with the highest curvature-penalized exploration score in the first region of the routing will be selected for high-speed trajectory planning. Then, inspired by [28], we propose an aggressive and safe exploration-oriented (ASEO) trajectory planning method for high-speed continuous exploration. The ASEO trajectory is composed of three parts: an exploring part, a continuous part, and a safety part. The time of the exploring part, the time of the continuous part, and the feasibility of the safety part are optimized simultaneously to compute an aggressive trajectory for fast exploration and a backup trajectory for safety.
@@
We evaluate the proposed method by comparing it against state-of-the-art methods in various challenging environments. The results show that the proposed method has $5.2\%-89.7\%$ higher exploration efficiency, $10.4\%-69.7\%$ faster UAV speed, and $28.9\%-93.9\%$ lower computational cost. Further, the proposed method is deployed on a UAV and is tested in real-world environments to evaluate its effectiveness. The contributions of this work are summarized as follows:
@@
    We propose an exploration-oriented heuristic double-tree (EOHDT) algorithm for efficiently solving large-scale ATSP. EOHDT computes a satisfactory region routing with real-time performance even in hundred-meter-level environments.
@@
    To achieve high-speed continuous exploration, we propose a curvature-penalized continuous score evaluation for the determination of short-term exploration targets.
@@
    We propose the ASEO trajectory planning method that ensures trajectory efficiency and safety simultaneously.
@@
    Extensive comparative simulations are conducted to evaluate the proposed method. We also test the proposed method by real-world experiments. We will release our source code.
@@ II-A Greedy Exploration Planning
The frontier-based method is first proposed by Yamauchi et al. [12]. It selects the frontier voxel that is closest to the robot as the exploration target. To minimize the UAV velocity change, the work in [13] selects the frontier voxel that is located at the boundary of the camera FOV as the exploration target. Different from the greedy frontier-based methods that only consider the next step exploration, Bircher et al. [14] proposes RH-NBVP, which samples a set of exploration motion sequences and executes the most informative sequence’s first step, and it shows high efficiency in short-term exploration. Nevertheless, in large-scale environments, it suffers from sample failure. To cope with this problem, Selin et al. [29] use frontiers to guide sampling when a local sample fails. However, long-distance sampling is time-consuming, and it does not guarantee success even with the guidance of the frontier. To reduce the planning cost and improve long-distance exploration, road graphs, which can compress the complex explorable space information and traversable space information into compact nodes and edges, are widely used in recent proposed greedy methods [15, 30, 17, 9, 31, 32, 33]. Consequently, greedy methods are able to plan efficiently on road graphs in large environments. Although these methods show high efficiency in short-term exploration and low computational cost, they suffer from back-and-forth maneuvers due to a lack of long-term consideration.
@@ II-B Optimal-Routing-Based Exploration Planning
Optimal-routing-based methods usually formulate the exploration target decision problems into routing problems. Optimal-tour-based methods can be roughly devided into frontier-ATSP-based methods [19, 34, 20] and coverage-ATSP-based methods [35, 22, 23, 6, 21]. Frontier-ATSP-based methods formulate ATSP using frontiers and find the shortest tour between frontiers. Coverage-ATSP-based methods formulate ATSP with unexplored space and find the shortest coverage path. The main drawback of the optimal-routing-based methods is that they have to spend an extremely high computational cost on solving ATSP in large-scale environments. Moreover, recent works [31, 36] report that optimal-routing-based methods usually do not have an obvious advantage in minimizing the exploration path distance compared to greedy methods. This is because optimal-routing-based methods usually make implicit optimal assumptions. Frontier-ATSP-based methods assume new frontiers will not arise after exploring a frontier. Coverage-ATSP-based methods assume all the unexplored space is traversable for the robot. Obviously, the above assumptions do not hold in most cases. Therefore, spending too much effort on computing a high-quality ATSP solution is not cost-effective.
II-C Learning-Based Exploration Planning
Learning-based methods have been used to tackle autonomous exploration problems with the advancement of machine learning. The work in [37] presents a modular deep-reinforcement-learning-based exploration framework that uses a convolutional Q-network to select goals from partial maps. The works in [38, 27] employ modular reinforcement learning frameworks for efficient visual exploration in unknown environments and provide systematic evaluations across realistic settings. The works in [26, 25] propose to construct a graph representation of the environment as the input to a graph neural network. However, they rely on discrete action spaces sampled from grid maps, which limits the optimality of the final path. In addition, the loading capacities of UAVs are usually limited, making them hard to carry extra devices for parallel computing.
II-D Exploration-Oriented Trajectory Planning
The exploration trajectory plays a crucial role in exploration tasks. The early proposed exploration methods [14, 12] use stop-and-go motions. To utilize the good maneuverabilities of UAVs, Dharmadhikari et al. [18] adopts motion primitives to speed up the exploration. To further improve the trajectory speed, Zhou et al. [19] optimizes a minimum-time trajectory with optimizable terminal velocity and acceleration for high-speed exploration. Although the minimum-time trajectory is very efficient, its optimizable terminal velocity and acceleration do not ensure improving the exploration continuity when new frontiers are generated after exploring the old frontier. What is more, the optimizable terminal velocity and acceleration can lead to a collision without considering obstacles that may be detected in the unknown space.
@@ III SYSTEM OVERVIEW
The dual-layer exploration target decision module plans both the long-term region routing and the short-term viewpoint targets. The trajectory planning module is responsible for planning high-speed continuous exploration trajectories. To achieve fast decision-making, a dynamic voxel map with resolution $r_{v}$ is updated by integrating depth image and odometry information. Different from the normal voxel map, the dynamic voxel map only maintains a local part of voxels around the UAV, while other voxels are dynamically read and written to save memory usage in large environments. Based on the voxel map, a dynamic topological graph (DTG, Fig. 2(a)) is maintained to manage exploration-related information. The traversable information is maintained by DTG’s history nodes and edges between them, and the explorable space information is maintained by DTG’s explorable region of interest (EROI), which is a set of uniform-scale cubic regions that can be explored by visiting corresponding active viewpoints. For detailed information, we refer readers to our previous work [39]. The exploration planning utilizes DTG to efficiently formulate ATSP, and ATSP is solved through EOHDT to find satisfactory exploration region routing with small computation. Then, three viewpoints, i.e., an exploring viewpoint, a continuous viewpoint, and a safety viewpoint in the first region of the region routing, are selected for the following trajectory planning. With these three viewpoints, an aggressive and safe exploration-oriented trajectory is generated, which is divided into three parts from the exploring viewpoint. The first part is named the exploring part, which goes from the UAV’s current position to the exploring viewpoint. The second part is named the continuous part, which goes from the exploring viewpoint to the continuous viewpoint. The last part is named the safety part, which goes from the exploring viewpoint to the safety viewpoint. After every replanning, the exploring part of the aggressive and safe exploration-oriented trajectory will be actually executed.
@@ IV DUAL-LAYER EXPLORATION VIEWPOINTS PLANNING
In this section, we first propose the EOHDT algorithm for finding a satisfactory region routing efficiently in large-scale environments. Then, we use a curvature-penalized continuous score to decide the targets in the first region of the tour for the following high-speed ASEO trajectory planning.
@@
Figure 2: An illustration of DTG and an instance of EOHDT-based routing planning. (a) DTG and its boundary regions (pink dashed boxes). (b) An MST that connects all the boundary regions of DTG is extracted. (c) The leaf point, whose distance is the longest to the root node, and its corresponding branch (green) will be executed last. The numbers indicate the lengths of edges. (d) Once there are multiple child nodes for the current node, the child node (orange dashed curve) whose first node has the shortest distance to the current node will be chosen as the next branch. (e) The final routing found by the proposed EOHDT-based routing planning.
IV-A EOHDT-Based Global Routing Planning
@@
As analyzed in Section II, a far-sighted and computationally efficient exploration planner for the large-scale exploration tasks is required. Therefore, instead of finding a high-quality routing of all precise explorable frontiers, the proposed method aims to spend low computation on finding a satisfactory coarse routing of boundary regions. We propose the EOHDT-based routing planning, which has a low computational complexity and a length guarantee in the worst case, for solving ATSP. The process of EOHDT-based routing planning is shown in Fig. 2(b-e).
@@
Firstly, we extract all the boundary regions. Here, a boundary region is composed of a history node and a set of EROIs (at least one) connected to this history node as shown in Fig. 2(a). To reduce the computation when formulating ATSP, the distances between all boundary regions and the distances between the robot and all boundary regions are computed by conducting the Dijkstra algorithm on DTG. Thanks to the sparse DTG, the ATSP can be formulated efficiently.
@@
To solve ATSP efficiently, we design an EOHDT algorithm, which is improved from the double-tree algorithm [40] and tailored to ATSP. The key idea of the double-tree algorithm is to construct a minimum spanning tree (MST, Fig. 2(b)) using the Prim algorithm and find a walk that goes around the MST. The solution routing of the double-tree algorithm ensures no longer than twice the optimal routing. However, the double-tree algorithm is designed for symmetric TSP. For exploration-oriented ATSP, the robot does not have to return to the original node as stated by [19]. Therefore, to minimize the routing length, the leaf node that is the farthest from the robot and nodes on its corresponding branch (in Fig. 2(c), the green nodes) will be executed last to save the highest cost of returning to the original node. Another problem is how to decide the sequence of the other branches. Since all the regions will be explored finally, the proposed method chooses the next node closest to the current node heuristically when the current node has multiple child nodes, as shown in Fig. 2(d). Finally, a routing for ATSP will be found as shown in Fig. 2(e).
Theorem 1.
The EOHDT runs in $\mathcal{O}(n^{2})$ time, where $n$ denotes the number of the history nodes.
Proof.
EOHDT starts by generating an MST utilizing the Prim algorithm, which takes $\mathcal{O}(n^{2})$ time. Then, EOHDT finds the farthest leaf node among $n_{leaf}$ leaf nodes, which takes $\mathcal{O}(n_{leaf})$ time. Let $m$ be the number of nodes that have multiple child nodes in the MST. Then, EOHDT needs to conduct $m$ times sorting of child nodes. It results in $\mathcal{O}(\sum_{i=1}^{m}n_{i}\log n_{i})$ time complexity, where $n_{i}$ is the child nodes number of the $i$-th node among the $m$ nodes. Because $n_{i}<n$ and $2\leq n_{i}$, we have $\sum_{i=1}^{m}n_{i}\log n_{i}<n^{2}$. Therefore, EOHDT runs in $\mathcal{O}(n^{2}+n_{leaf}+\sum_{i=1}^{m}n_{i}\log n_{i})\approx\mathcal{O}(n^{2})$ time.
Theorem 2.
@@
The length of the routing solved by EOHDT $l_{E}$ holds $l_{E}\leq 2l^{*}-l_{m}$, where $l^{*}$ is the length of the optimal solution of the ATSP, and $l_{m}$ is the longest distance among distances between the robot and all the nodes on the MST.
Proof.
Given the summation of all the edge lengths of the MST $l_{MST}$, it holds $l_{MST}\leq l^{*}$. It is easy to derive
${\color[rgb]{0.0,0.0,0.0}l_{E}=2l_{MST}-l_{m}\leq 2l^{*}-l_{m}.}$ (1)